Problem 9-25 Direct Materials and Direct Labor Variances; Computations from Incomplete Data [LO9-4, LO9-5]

Sharp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
Direct materials 3 feet $ 5 per foot $ 15
Direct labor ? hours ? per hour ?

During March, the company purchased direct materials at a cost of $55,650, all of which were used in the production of 3,200 units of product. In addition, 4,900 direct labor-hours were worked on the product during the month. The cost of this labor time was $36,750. The following variances have been computed for the month:

Materials quantity variance $ 4,500 U
Labor spending variance $ 3,150

U

Labor efficiency variance $ 700

U

Required:

1. For direct materials:

a. Compute the actual cost per foot of materials for March.

b. Compute the price variance and the spending variance.

2. For direct labor:

a. Compute the standard direct labor rate per hour.

b. Compute the standard hours allowed for the month’s production.

c. Compute the standard hours allowed per unit of product.
Answer #1

1a) Material quantity variance = (Standard qty-actual qty)Standard price

-4500 = (3200*3*5-5X)

5X = 52500

X(actual qty) = 10500

Actual cost per foot = 55650/10500 = 5.30

1b) Material price variance = (5-5.30)*10500 = 3150 U

Material spending variance = (3200*15-55650) = 7650 U

2a) Labor rate variance = (Standard rate-actual rate)actual hour

-2450 = (X*4900-36750)

-4900X = -34300

X(Standard rate) = 7

2b) Labor efficiency variance = (Standard hour-actual hour)Standard rate

-700 = (X*7-4900*7)

-7X = -33600

X(Standard hour) = 4800 Hour

2c) Standard hour per unit = 4800/3200 = 1.50 hour per unit